A block of with a mass of 4.0 kg sitting in a frictionless horizontal table is connected to a hanging block of mass 6.0 kg by a string that passes over a pulley. Assuming that friction, the mass of the string, and the mass of the pulley are negligible, at what rate do the blocks accelerate.


1
Expert's answer
2021-06-21T11:38:54-0400
"(m+M)a=Mg\\(4+6)a=6(9
.8)\\a=4.3\frac{m}{s^2}"